OpenStudy (anonymous):

The limit of √x-4 as x approaches 4

OpenStudy (anonymous):

careful here

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

note that while \(\sqrt{4-4}=\sqrt{0}=0\) the limit means the limit from both directions, right and left so while \(\lim_{x\to 4^+}\sqrt{x-4}=0\) is good, you cannot approach \(4\) from numbers below \(4\) becuase if \(x<4\) then \(x-4<0\) and so \(\sqrt{x-4}\) is not a real number

OpenStudy (anonymous):

what i am trying to say is that \[\lim_{x\to 4^-}\sqrt{x-4}\] does not exist, because the expression \(\sqrt{x-4}\) is only defined for \(x\geq 4\)

OpenStudy (anonymous):

Ok that makes a lot of sense. Thank you!

OpenStudy (anonymous):

yw